I could really use some help with this problom, please. I can't make the answer come out as an inequality. Thank-you
-19<=3x-6<=-5

Answer by edjones(7311)

-19<=3x-6<=-5
.
-19<=3x-6
-13<=3x
-13/3<=x
.
3x-6<=-5
3x<=1
x<=1/3
.
-13/3<=x<=1/3 or x[-13/3, 1/3]
.
Ed